# Towards Aligning Language Models with Textual Feedback

### v2

@@ 3.1 Toxicity Reduction
We follow the same experimental setup as in Liu et al. (2021); Lu et al. (2022), and consider reducing toxicity from GPT2-large. As a reward function and a proxy for measuring the toxicity of the LLM generations, we use the Perspective API. [^3] The Perspective API is a service developed by Google that is dynamic and evolves over time. Queries were made from Sep 2023 to Nov 2023. We use $K=5$ quantiles, obtained by sorting the samples in the data pool from lowest toxicity (highest reward) to highest toxicity (lowest reward), and map them to language feedback indicating increasing degrees of toxicity: ‘‘Lowest Toxicity’’, ‘‘Low-Moderate Toxicity’’, ‘‘Moderate Toxicity’’, ‘‘High-Moderate Toxicity’’, and ‘‘Maximum Toxicity’’. We report training details and hyper-parameters in A.1.
@@
During evaluation, we sample 25 generations for each prompt using nucleus sampling with $top\_p=0.9$ and condition on Lowest Toxicity. We report the avg. max. toxicity, measured as the average maximum toxicity over the 25 generations, and the toxicity prob. as the empirical toxic probability of at least one of any 25 generations being toxic, i.e., $score>0.5$ (both measured by PerspectiveAPI). Regarding language quality, the fluency is measured as the conditional output perplexity of a response given a prompt according to a larger GPT2-XL model, which acts as a proxy for how much our trained model deviates from the initial policy. We also compute diversity as the number of distinct n-grams normalized by the total length of the text.
@@ 3.2 Summarization
We next evaluate if ALT can better help align the LLM to user preferences. To verify this, we experiment using the Reddit TL;DR dataset Völske et al. (2017). The TL;DR dataset has become a common benchmark for measuring alignment. We used the dataset version hosted in HuggingFace [^4] https://huggingface.co/datasets/CarperAI/openai_summarize_tldr, with splits amounting to 117k, 6.5k, and 6.5k for train, validation, and test. The prompts consist of Reddit forum posts and the task is to generate a summary of the main points of the post while fulfilling different facets that humans care about, such as coherence, accuracy, coverage, or conciseness. Once again, we use an existing reward model that accounts for the different facets humans value on summaries to predict a single scalar reward and quantize the reward model’s feedback into preset texts indicating increasing degrees of alignment fulfillment.
@@
We implement a version of Quark, with a slight modification as to sample multiple generations per prompt to compute the reward quantiles locally instead of globally across all prompts. We found that this was crucial for training. We use $K=5$ quantile tokens, which are newly added to the tokenizer. We sample 96 generations for each prompt but only train on 10 generations drawn at random (2 for each quantile) to speed up training. On top of that, $\textsc{ALT}_{\textsc{RM}}$ is implemented by mapping reward quantiles to textual feedback. We prepend to the prompt the feedback sentence formatted as ‘‘<feedback > input: ’’; where the language feedback is one of: ‘‘Excellent’’, ‘‘Good’’, ‘‘Mediocre’’, ‘‘Bad’’, and ‘‘Horrible’’. Similarly, 96 generations per prompt are sampled though training takes place only on 10 samples (2 for each feedback type).
@@ 4 Results
Compared to SteerLM, we find that $\textsc{ALT}_{\textsc{LM}}$ is better off at learning to discriminate between ‘‘Harmless and very helpful’’ and ‘‘Harmless and helpful’’, as it almost monotonically increases the % of the former and decreases the % of the latter. We observe that SteerLM fails at improving both the fraction of generations being ‘‘Harmless and very helpful’’ and ‘‘Harmless and helpful’’ compared to the SFT model, and that it becomes less harmful by becoming less helpful, e.g., denying to answer a question or providing an unrelated answer. Regarding ‘‘Harmless and not helpful’’ and ‘‘Harmful’’, both methods follow similar trends but $\textsc{ALT}_{\textsc{LMC}}$ achieves better performance. We include other training curves showing the generations’ length, percentage of truncated generations, and perplexity in A.3, and also perform preliminary tests on $\textsc{ALT}_{\textsc{LMC}}$’s ability to steer the LM in Appendix E. We provide qualitative results in B.3.
@@ A.1 Toxicity Reduction
We fine-tune GPT2-large using the following language feedback tags: "Lowest Toxicity", "Low-Moderate Toxicity", "Moderate Toxicity", "High-Moderate Toxicity", and "Maximum Toxicity". At inference time we target nontoxic behavior by sampling conditioned on the "best" feedback type, i.e., "Lowest Toxicity". As each element in the batch might have associated language feedback tokens of different lengths, we pad them on the left to match the size of the longest feedback tokens within the current batch. We also insert a newly added separator token “<|separator>|” between the feedback tokens and the query input IDs, which is useful for easy removal of the feedback tokens when required on different points of the training pipeline.
@@
In this experiment, the KL-penalty term with the reference policy in the loss function was important in avoiding obtaining a low-toxicity policy that would just output gibberish language. However, in the subsequent experiments, we got rid of this KL-penalty term without sacrificing perplexity, reducing thus the need for storing the reference policy during training. We hypothesize that for the unlearning toxicity task, this was needed as we departed training from a pre-trained model and because the task was to complete text from a few query tokens, as opposed to starting from an SFT model and having a more clearly defined task on summarization and dialogue.
@@ A.2 Summarization
We fine-tune the GPT-J SFT model using the language feedback mentioned in subsection 3.2. As the distinct phases of our algorithm are decoupled, one can leverage different computation resources at every stage. The data collection stage is the most costly one in terms of time required to sample and provide feedback to several generations, but one can launch multiple smaller GPU jobs and employ frameworks for faster inference such as vLLM Kwon et al. (2023) for substantial speedup. We carried out the sampling phase on 8 parallel NVIDIA 3090 (24GB) threads that sampled on different disjoint subsets of the 2048 prompts. Following this setup, the $2048*96=196608$ generations can be sampled in less than 10min. The same can be applied for the feedback stage, either for Quark and $\textsc{ALT}_{\textsc{RM}}$, which only require running the Reward Model and can be done in several independent NVIDIA 3090 GPUs; or for $\textsc{ALT}_{\textsc{LM}}$ in which several independent CPU-only jobs can query the GPT-3.5 model through the OpenAI API. For all experiments, training is done in 2 NVIDIA A100 (80GB) and employing DeepSpeed Rasley et al. (2020) for handling the training parallelization and leveraging the ZeRO-optimizer and CPU offloading features. Each training iteration takes  3-4h to complete.
@@ A.3 Dialogue
We fine-tune the Pythia-2.8b SFT model using the language feedback mentioned in subsection 3.3. The same decoupled nature for the data collection and training described in A.2 applies here, and we leveraged the same computation resources. Figure 5 contains the generations’ length and % of truncated generations along iterations for $\textsc{ALT}_{\textsc{LMC}}$ and SteerLM, and Figure 6 contains the perplexity curve as training progresses.
@@ B.2 Summarization
We include qualitative example summaries generated by SFT, PPO, Quark, $\textsc{ALT}_{\textsc{RM}}$, and $\textsc{ALT}_{\textsc{LMU}}$ policies for the TL;DR Reddit posts in Figure 7, Figure 8, Figure 9, Figure 10, Figure 11, Figure 12, and Figure 13.
@@ B.3 Dialogue
We include qualitative examples generated by SFT, DPO, SteerLM, and $\textsc{ALT}_{\textsc{LMC}}$ policies for the Anthropic-HH single-turn user queries in Figure 14, Figure 15, Figure 16, and Figure 17.
In Figure 18, we also include some failed examples for which we observed that the GPT-3.5-Turbo feedback is not aligned with what a user would prefer. Due to the limited coverage of the in-context examples on the GPT-3.5-Turbo prompt, the model might sometimes fail at providing feedback representative of human preferences. For instance, GPT-3.5-Turbo might consider a generation to be very helpful when the Assistant states that it will be glad to help the user (but no actual helpful response), or when the Assistant states that it includes a video, or when it provides a verbose non-factual response. Similarly, GPT-3.5-Turbo may consider a generation to be harmful influenced by the harmfulness of the user query even if the Assistant dodges the request.
We believe that, as every Reward Model, GPT-3.5-Turbo can be noisy and faulty. In our results in Figure 3, we showed that we can steer a policy towards producing more helpful and harmless responses as perceived by GPT-3.5-Turbo, which might not be sufficient for user-facing applications. However, we expect that improving the capacity of the feedback provider model and ensuring that it is aligned with human standards of quality would lead to a better policy model.
Figure 14: Qualitative results example 1 on HH-dialogue. Chosen refers to the annotator’s chosen response over a pair of responses from the Anthropic-HH dataset and it’s an off-policy generation. Notice that DPO did not train on the GPT-3.5-Turbo feedback, but directly on the Anthropic-HH preference dataset.
Figure 15: Qualitative results example 2 on HH-dialogue. Chosen refers to the annotator’s chosen response over a pair of responses from the Anthropic-HH dataset and it’s an off-policy generation. Notice that DPO did not train on the GPT-3.5-Turbo feedback, but directly on the Anthropic-HH preference dataset.
Figure 16: Qualitative results example 3 on HH-dialogue. Chosen refers to the annotator’s chosen response over a pair of responses from the Anthropic-HH dataset and it’s an off-policy generation. Notice that DPO did not train on the GPT-3.5-Turbo feedback, but directly on the Anthropic-HH preference dataset.
Figure 17: Qualitative results example 4 on HH-dialogue. Chosen refers to the annotator’s chosen response over a pair of responses from the Anthropic-HH dataset and it’s an off-policy generation. Notice that DPO did not train on the GPT-3.5-Turbo feedback, but directly on the Anthropic-HH preference dataset.
Figure 18: Examples on noisy/faulty GPT-3.5-Turbo feedback. Top: GPT-3.5-Turbo wrongly identifies a response as very helpful when the Assistant states that it has a recipe, though no recipe is provided. Middle: GPT-3.5-Turbo wrongly assesses a verbose, non-accurate response as being very helpful. Bottom: GPT-3.5-Turbo wrongly judges a response as harmful, influenced by the harmfulness of the Human query.
@@
Figure 19: GPT-3.5 few-shot prompt for providing categorical feedback on $\textsc{ALT}_{\textsc{LMC}}$ for the HH dialogue task.
@@ Appendix D SteerLM implementation details
Appendix E HH test-time steerability assessment
We conducted a small evaluation setting to assess the steerability of our model at run-time. To do so, we evaluated our last model checkpoint when conditioned on Harmful instead of Harmless and very helpful. We notice that this leads to generations being slightly less Harmless and very helpful (65.1 $\rightarrow$ 60.0) and slightly more Harmful (11.6 $\rightarrow$ 15.1), but that our model is not steerable at inference time to produce harmful content. This is consistent with our exploitation of the exemplar feedback commented in 2.3.1. Similarly to PPO and DPO, by using a single exemplar feedback, we are steering our model towards an implicit set of user preferences, i.e., being harmless and very helpful, and thus our model struggles when asked to elicit the opposite behavior.
Appendix F Examples of unconstrained feedback for $\textsc{ALT}_{\textsc{LMU}}$ on TL;DR

### v3

@@ 3.1 Toxicity Reduction
We follow the same experimental setup as in Liu et al. (2021); Lu et al. (2022), and consider reducing toxicity from GPT2-large. As a reward function and a proxy for measuring the toxicity of the LLM generations, we use the Perspective API. [^3] The Perspective API is a service developed by Google that is dynamic and evolves. Queries were made from Sep 2023 to Nov 2023. We use $K=5$ quantiles, obtained by sorting the samples in the data pool from lowest toxicity (highest reward) to highest toxicity (lowest reward), and map them to language feedback indicating increasing degrees of toxicity: ‘‘very toxic’’, ‘‘toxic’’, ‘‘medium toxic’’, ‘‘slightly toxic’’, ‘‘nontoxic’’. We report training details and hyper-parameters in A.1.
@@
During evaluation, we sample 25 generations for each prompt using nucleus sampling with $top\_p=0.9$ and condition on nontoxic. We report the avg. max. toxicity, measured as the average maximum toxicity over the 25 generations, and the toxicity prob. as the empirical toxic probability of at least one of any 25 generations being toxic, i.e., $score>0.5$ (both measured by PerspectiveAPI). Regarding language quality, the fluency is measured as the conditional output perplexity of a response given a prompt according to a larger GPT2-XL model, which acts as a proxy for how much our trained model deviates from the initial policy. We also compute diversity as the number of distinct n-grams normalized by the total length of the text.
@@ 3.2 Summarization
We next evaluate if ALT can better help align the LLM to user preferences. We experiment using the Reddit TL;DR dataset Völske et al. (2017) to verify this. The TL;DR dataset has become a common benchmark for measuring alignment. We used the dataset version hosted in HuggingFace [^4] https://huggingface.co/datasets/CarperAI/openai_summarize_tldr, with splits amounting to 117k, 6.5k, and 6.5k for train, validation, and test. The prompts consist of Reddit forum posts and the task is to generate a summary of the main points of the post while fulfilling different facets that humans care about, such as coherence, accuracy, coverage, or conciseness. Once again, we use an existing reward model that accounts for the different facets humans value on summaries to predict a single scalar reward and quantize the reward model’s feedback into preset texts indicating increasing degrees of alignment fulfillment.
@@
We implement a version of Quark, with a slight modification as to sample multiple generations per prompt to compute the reward quantiles locally instead of globally across all prompts. We found that this was crucial for training. We use $K=5$ quantile tokens, which were recently added to the tokenizer. To speed up training, we sample 96 generations for each prompt but only train on 10 generations drawn at random (2 for each quantile). On top of that, $\textsc{ALT}_{\textsc{RM}}$ is implemented by mapping reward quantiles to textual feedback. We prepend to the prompt the feedback sentence formatted as ‘‘<feedback > input: ’’; where the language feedback is one of: ‘‘Excellent’’, ‘‘Good’’, ‘‘Mediocre’’, ‘‘Bad’’, and ‘‘Horrible’’. Similarly, 96 generations per prompt are sampled though training takes place only on 10 samples (2 for each feedback type).
@@ 4 Results
Compared to SteerLM, we find that $\textsc{ALT}_{\textsc{LM}}$ is better off at learning to discriminate between ‘‘Harmless and very helpful’’ and ‘‘Harmless and helpful’’, as it almost monotonically increases the % of the former and decreases the % of the latter. We observe that SteerLM fails at improving both the fraction of generations being ‘‘Harmless and very helpful’’ and ‘‘Harmless and helpful’’ compared to the SFT model, and that it becomes less harmful by becoming less helpful, e.g., denying to answer a question or providing an unrelated answer. Regarding ‘‘Harmless and not helpful’’ and ‘‘Harmful’’, both methods follow similar trends but $\textsc{ALT}_{\textsc{LMC}}$ achieves better performance. We include other training curves showing the generations’ length, percentage of truncated generations, and perplexity in A.3. We provide qualitative results in B.3.
@@ A.1 Toxicity Reduction
We fine-tune GPT2-large using the following language feedback tags: ‘‘very toxic’’, ‘‘toxic’’, ‘‘medium toxic’’, ‘‘slightly toxic’’, ‘‘nontoxic’’. At inference time we target nontoxic behavior by sampling conditioned on the "best" feedback type, i.e., "nontoxic". As each element in the batch might have associated language feedback tokens of different lengths, we pad them on the left to match the size of the longest feedback tokens within the current batch. We also insert a newly added separator token “<|separator>|” between the feedback tokens and the query input IDs, which is useful for easy removal of the feedback tokens when required on different points of the training pipeline.
@@
In this experiment, the KL-penalty term with the reference policy in the loss function was important in avoiding obtaining a low-toxicity policy that would just output gibberish language. However, in the subsequent experiments, we got rid of this KL-penalty term without sacrificing perplexity, thus reducing the need for storing the reference policy during training. We hypothesize that for the unlearning toxicity task, this was needed as we departed training from a pre-trained model and because the task was to complete text from a few query tokens, as opposed to starting from an SFT model and having a more clearly defined task on summarization and dialogue.
@@ A.2 Summarization
We fine-tune the GPT-J SFT model using the language feedback mentioned in subsection 3.2. As the distinct phases of our algorithm are decoupled, one can use different computation resources at every stage. The data collection stage is the most costly one in terms of time required to sample and provide feedback to several generations, but one can launch multiple smaller GPU jobs and employ frameworks for faster inference such as vLLM Kwon et al. (2023) for substantial speedup. We carried out the sampling phase on 8 parallel NVIDIA 3090 (24GB) threads that sampled on different disjoint subsets of the 2048 prompts. Following this setup, the $2048*96=196608$ generations can be sampled in around 10min. The same can be applied for the feedback stage, either for Quark and $\textsc{ALT}_{\textsc{RM}}$, which only require running the Reward Model and can be done in several independent NVIDIA 3090 GPUs; or for $\textsc{ALT}_{\textsc{LM}}$ in which several independent CPU-only jobs can query the GPT-3.5 model through the OpenAI API. For all experiments, training is done using 2 NVIDIA A100 (80GB) and employing DeepSpeed Rasley et al. (2020) for handling the training parallelization and leveraging the ZeRO-optimizer and CPU offloading features. Each training iteration takes  3-4h to complete.
@@ A.3 Dialogue
We fine-tune the Pythia-2.8b SFT model using the language feedback mentioned in subsection 3.3. The same decoupled nature for the data collection and training described in A.2 applies here, and we used the same computation resources. Figure 5 contains the generations’ length and % of truncated generations along iterations for $\textsc{ALT}_{\textsc{LMC}}$ and SteerLM, and Figure 6 contains the perplexity curve as training progresses.
@@ B.2 Summarization
We include qualitative example summaries generated by SFT, PPO, Quark, $\textsc{ALT}_{\textsc{RM}}$, and $\textsc{ALT}_{\textsc{LMU}}$ policies for the TL;DR Reddit posts in Figure 7, and Figure 8.
@@ B.3 Dialogue
We include qualitative examples generated by SFT, DPO, SteerLM, and $\textsc{ALT}_{\textsc{LMC}}$ policies for the Anthropic-HH single-turn user queries in Figure 9.
Figure 9: Qualitative results example 1 on HH-dialogue. Chosen refers to the annotator’s chosen response over a pair of responses from the Anthropic-HH dataset.
@@
Figure 10: WARNING: Contains harmful examples from the HH dataset GPT-3.5 few-shot prompt for providing categorical feedback on $\textsc{ALT}_{\textsc{LMC}}$ for the HH dialogue task.
@@ Appendix D SteerLM implementation details
Appendix E Examples of unconstrained feedback for $\textsc{ALT}_{\textsc{LMU}}$ on TL;DR